# Molecular Learning Dynamics

### v1

@@ Abstract
We apply the physics–learning duality to molecular systems by complementing the physical description of interacting particles with a dual learning description, where each particle is modeled as an agent minimizing a loss function. In the traditional physics framework, the equations of motion are derived from the Lagrangian function, while in the learning framework, the same equations emerge from learning dynamics driven by the agent loss function. The loss function depends on scalar quantities that describe invariant properties of all other agents or particles. To demonstrate this approach, we first infer the loss functions of oxygen and hydrogen directly from a dataset generated by the CP2K physics-based simulation of water molecules. We then employ the loss functions to develop a learning-based simulation of water molecules, which achieves comparable accuracy while being significantly more computationally efficient than standard physics-based simulations.
@@ 1 Introduction
In this article, we propose a novel framework for modeling molecular systems by applying the physics-learning duality [14], where the physics description of interacting nuclei is complemented by a machine learning description of interacting agents. [^1] A complementary approach was taken in Ref. [15] where a machine learning view of autonomous agents was replaced with a dual physical view of autonomous particles. The agents (or nuclei) navigate in a complex environment of other agents (or nuclei), and the same physical equations of motion emerge from learning dynamics governed by agent loss functions. These loss functions depend on scalar quantities that describe invariant properties of other agents, allowing the system to retain key symmetries at the architectural level. While the loss function could in principle be derived from first principles, in this article we infer it directly from the CP2K simulations [16, 17] of water molecules. We then use the learned loss function to construct a learning-based simulation of water molecules, which runs many orders of magnitude faster than conventional physics-based molecular dynamics algorithms.

### v2

@@ Abstract
We apply the physics-learning duality to molecular systems by complementing the physical description of interacting particles with a dual learning description, where each particle is modeled as an agent minimizing a loss function. In the traditional physics framework, the equations of motion are derived from the Lagrangian function, while in the learning framework, the same equations emerge from learning dynamics driven by the agent loss function. The loss function depends on scalar quantities that describe invariant properties of all other agents or particles. To demonstrate this approach, we first infer the loss functions of oxygen and hydrogen directly from a dataset generated by the CP2K physics-based simulation of water molecules. We then employ the loss functions to develop a learning-based simulation of water molecules, which achieves comparable accuracy while being significantly more computationally efficient than standard physics-based simulations.
@@ 1 Introduction
In this article, we propose a novel framework for modeling molecular systems by applying the physics-learning duality [14, 15], where the physics description of interacting nuclei is complemented by a machine learning description of interacting agents. [^1] A complementary approach was taken in Ref. [16] where a machine learning view of autonomous agents was replaced with a dual physical view of autonomous particles. The agents (or nuclei) navigate in a complex environment of other agents (or nuclei), and the same physical equations of motion emerge from learning dynamics governed by agent loss functions. These loss functions depend on scalar quantities that describe invariant properties of other agents, allowing the system to retain key symmetries at the architectural level. While the loss function could in principle be derived from first principles, in this article we infer it directly from the CP2K simulations [17, 18] of water molecules. We then use the learned loss function to construct a learning-based simulation of water molecules, which runs many orders of magnitude faster than conventional physics-based molecular dynamics algorithms.